OpenStudy (anonymous):

A sample of 143 firecrackers contained 5 duds. How many duds would you expect in a sample of 1430 firecrackers?

OpenStudy (anonymous):

10 times 5

OpenStudy (anonymous):

143/5 = 1430/x 143x=7150 x=50 duds

OpenStudy (anonymous):

50